Skip to Search

 
diagonal spacer

Information Technology Staff Meeting Notes
March 27, 2009

Present: Thomas Bielicke, Cynthia Bowker, Laurie Brennan, Will Chaney, Logan Clapp, Debra Fetzer, Jason Fleming, Winston Harris, Jay, Bobbie Meng, Tom Minton, Chris Nicolich, Amy Polk, Patricia Ruwell, Mark Sullivan, Evan Wack, Stephen Williams
Announcements
  1. This is Will's last staff meeting as interim head of IT. Bill returns next week.
  2. Judy has rescheduled her attendance at this meeting to Friday, April 3.
  3. Employee Recognition Meeting celebration - next Friday is the Spring Picnic - need to reschedule.
  4. Matthew's wife had the baby - a boy.
Ongoing IT Updates
Equipment Received/Installed/Repaired
  1. Orders - SAN expansion - filling vacant drive space, plus 10 drives for DLC growth - possibly including PALMM.
  2. IT action plan - server center on Waldo Road - consolidation off site.
  3. Logan closed down all the 4400's yesterday.
  4. Cynthia notes backups are close on space.
  5. Workstation migration - done. Hesham's machine in Documents is still on Smatherslib. Joe wants to check some documents, and then it can be wiped.
  6. Laptop encryption - laptop 14 is getting an "invalid recovery error".
Current Projects
  1. Admin rights for DLC - created a policy group. What about room 1A users? It can be fine tuned for USB users. Need to create a ticket.
  2. JM-ED VLAN changes - done. Xerox has not been updated in JM. Should be done next week. Education public printing is back. CNS can wait one more week for JM.
  3. Auto play - on hold.
  4. SCCM to UFAD - it was ready to install on the server this morning, but there was a hardware issue and we were unable to roll it out. We will try it on VM. Remote deployment for server OS - VISTA & 2K8.
  5. EPO - Ariel & print daemons machines are working fine with the new policy.
  6. Green Team monitor request - April 20.
Hardware Support Unit
  1. Inventory - IT finished - vaults, etc. Evan will take the van full of items to be surveyed to Property today. Very close to completely done with inventory.
  2. Updates - SUS - working on all public machines except West. Logan would like it to be noted whether DHCP or static. Amy wants non-pings reported via Grover ticket.
  3. Testing Steady State on Reference/Circ desks next week.
  4. Do we support Chrome? Yes.
  5. UFIRT notification - Peter McKay's database machine on public table 10 in West - CNS says it must be rebuilt. Evan says he scanned it. IP 193.42
  6. NT16 - out of disk space. No list of tapes in the directory. It was offline this morning. It said it was in use, but there was no job running. Will look at NT16 & NT18 backups this afternoon.
Software Projects Status
  1. ARL stats - 33.3 percent complete. Milestone of one-third complete is about a week away.
  2. Root cert - THAWTE. Need to move on it.
  3. NT4 - Chris has moved all shortcut links off. Just need to move the configuration management folder and it will be complete.
  4. Logan made a wonderful set up for SQL upgrade. Just need to get the license. Need to figure out the CALS, and then put in the order.
  5. New development test server - tried to install Shibboleth - getting closer.
  6. ILLIAD taskforce to meet Monday. Migration moved to the break between spring and summer.
  7. Demoed Degree Info application to Bonnie this morning. We are taking in her feedback.
  8. Patron Update - incorporating barcode info - Chris is working on it.
  9. OCR toolkit automation - design approved by Laurie Taylor.
  10. Gator Scholar - Val is working on the layout.
  11. UFDC - Mark - new version has Google display of search results.
ALEPH
  1. Tuesday - report upgrade - Education had some issues.
  2. Production upgrade this Sunday.
  3. Daniel Cromwell - brought up via email - that the print.ini & alephcom.ini files were not overwritten in the upgrade. In the future, we may want to write into the files what we don't want overwritten. It may be a moot point since we are upgrading to version 19.
DLC
  1. JPG2K3 needs more drives.
  2. Mark Sullivan will be gone May 11 - June 22. They don't really want to change things until he come back.
  3. Most of the DLC stuff is on hold until more space is acquired.
Web Projects
  1. New A-Z site index. Link on the bottom of the home page.
  2. Music - headers moved. Background broken. Tom fixed.
  3. Visitors page was put into the Giving site. Barbara Hood does not want it there. Will redirect.
IT Security Management
  1. Meeting held.

 

University of Florida Home Page